The cost of a bus excursion was $180. This amount was to have been shared equally by all those who said they wished to go. However, 6 people failed to show up so that those who did go each had to pay $1.50 more. How many people actually went on the trip?
I just want to know how to get started I do not need the actual answer.

Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

let x people decide to go initially
..
cost per person = 180/x
...
6 persons opt out . number which went = x-6
...
so amount to be paid = 180 /(x-6)
...
180/(x-6) -180/x = 1.5

Answer by stanbon(75887)

The cost of a bus excursion was $180. This amount was to have been shared equally by all those who said they wished to go.
---
Let that number of people be "x".
The average cost per person is 180/x dollars.
------------
However, 6 people failed to show up so that those who did go each had to pay $1.50 more.
Now the number of people is "x-6".
Now the average cost per person is 180/(x-6) dollars.
--------------
How many people actually went on the trip?
---
Solve: new average - old average = 1.50 dollars
================
Cheers,
Stan H.